# From Cool Demos to Production-Ready FMware: Core Challenges and a Technology Roadmap

### v1

@@ IV-K Cross-cutting Issues that affect productionizing FMware
Lack of Automated Testing Capabilities. The lack of automated testing in FMware hinders efficient issue detection and continuous integration. Existing automated tools often miss the nuanced complexities in FMware, making them inadequate for production systems where manual testing is impractical due to scale. Without automation, development slows, increasing costs and the risk of undetected regressions [92, 1].
@@
Lack of Assertion-Based Unit Tests. Writing assertion-based unit tests for FMware is difficult due to the non-deterministic nature of FMs, making expected outputs hard to define [92, 77]. Traditional tests rely on fixed outcomes, but FMware’s varying outputs complicate this, especially in agent construction and memory management [77]. Despite these challenges, such tests are critical in production environments to detect regressions during refactoring or updates [92].
@@
Text-Based Evaluation Leads to Overestimation of Quality. Text-based evaluation in FMware often leads to an overestimation of quality by failing to capture deeper issues like relevance, accuracy, or alignment with business rules. Gao et al. [93] highlight that current text-based Natural Language Generation (NLG) evaluations often miss key aspects such as context and factual correctness. Similarly, Hasanbeig et al. [94] argue that relying solely on surface-level outputs can cause models to overfit to specific evaluation prompts, masking underlying performance issues.
@@
Ineffective and Inefficient AI-as-Judge Technologies. AI-as-Judge systems face inefficiency and ineffectiveness due to high costs and inherent biases. Thakur et al. [95] highlight that models often struggle with hallucinations and inaccuracies, making their outputs unreliable in complex decision-making tasks. Additionally, the computational expense of using advanced models like GPT-4 makes them impractical for ongoing evaluations, especially when they fail to meet specific business needs [95, 96]. These limitations, combined with the need for custom evaluation pipelines, reduce the feasibility of deploying such systems in production FMware.
@@
Difficulty navigating the regulatory and legal compliance minefield. Legal compliance and licensing complexities permeate all stages of the FMware lifecycle. FMs must align with the legal requirements of their deployment regions; for example, using ChatGPT-4-based FMware in China may not be viable due to local regulations [97, 98]. Licensing FM data, including synthetic data, RLHF data, and user feedback, further complicates compliance. If a model generates synthetic data, the licensing of the data it was trained on affects the licensing of the generated data. Similarly, integrating user feedback (e.g., ratings or thumbs up/down) into FMware complicates matters if users later opt out, as removing their data from the FM and related components becomes difficult. Traditional Software Bills of Materials (SBOMs) ensure compliance for conventional software, but FMware’s complexity (encompassing data, models, and FM-generated data) goes beyond what traditional SBOMs can track.
V Challenges
@@
TABLE I: Challenges and Recurrent Issues in Developing Production-Ready FMware
@@
In this section, we present our thematic analysis results and outline the key challenges in productionizing FMware. It is important to note that the challenges presented in this version of the paper address only a subset of the identified recurrent issues, as shown in Table I. Future versions will expand on this by incorporating additional challenges that cover all the recurrent issues across the FMware lifecycle.
@@
V-A Testing
Overview. Poor testing practices present significant roadblocks in the transition from demo FMware to production-ready systems. As identified in Section IV, the lack of robust, assertion-based unit tests, the absence of automated testing, and the over-reliance on surface-level text-based evaluation [92, 77, 1] often lead to overestimation of FMware quality. In real-world production settings, they result in unreliable outputs, latent defects, and increased regression costs. To address these issues, FMware require innovative testing mechanisms that account for the unique complexities of FMs and their non-deterministic behavior.
@@ V-A Testing
Moreover, the role of AI-as-judge technologies in testing pipelines adds another layer of complexity. Current AI judges, particularly those using large FMs like GPT-4, exhibit several critical flaws in evaluating correctness. These technologies often favor outputs based on superficial attributes, such as formatting or the length of responses rather than factual accuracy or alignment with real-world constraints [95, 99]. For instance, research shows that simply reversing the order of candidate responses can dramatically alter the AI’s judgment, resulting in false positives or negatives. The failure to maintain consistency in judging introduces significant risks when scaling FMware for production environments. Research has also points to a misalignment between AI decisions and human expectations. For example, Koo et al. [100] calculate the Rank-Biased Overlap (RBO) score to measure the agreement between human preferences and model evaluations in ranking-generated texts across 16 different LLMs (RBO varies from 0 to 1 and higher values indicate higher agreement). The average RBO was 0.44, with GPT-4 scoring 0.47. Another major issue arises from the associated cost and latency with AI-as-judge systems. Relying on large models like GPT-4 is computationally expensive and not viable for long-term or large-scale production-grade testing [10].
@@
– Next-generation AI-judge framework. The development of such a framework is crucial for improving the reliability of evaluations. Unlike existing solutions, this framework would guide developers in crafting tailored prompts that align with specific business logic and domain constraints. These judges would be trained to focus not just on the superficial aspects of FM outputs (e.g., format, length) but on deeper attributes like factual accuracy, consistency with previous outputs, and compliance with application-specific requirements. To train the judges in a more structured, guided, flexible, evolvable, and cost-effective manner, we foresee the use of curriculum engineering (more details in Section V-F). As a result, we would also obtain a more lightweight and efficient model, further contributing to cost-effectiveness.
V-B Observability
Overview. Observability in FMware systems is crucial to ensure transparency, traceability, and reliability throughout their lifecycle. However, as outlined in Section IV, FMware presents unique challenges in observability due to its reliance on dynamic, non-deterministic components like FMs and autonomous agents. Recurrent issues, such as the complexity in determining the rationale behind system failures (i.e., whether stemming from limitations in prompts or FMs themselves) and the lack of FMware-native observability mechanisms [1], highlight the need for novel solutions. Unlike traditional software, FMware systems demand observability approaches that can capture both functional and cognitive aspects of AI agents and models.
Critical Analysis of the State-of-the-Practice. Current observability tools in FMware, while evolving, remain grounded in classical software observability practices. Tools like OpenLLMetry [101] and LangSmith [102] focus on low-level resource monitoring, capturing traces of FM calls, Vector DB interactions, and user prompts, along with metrics like latency and resource utilization. However, these frameworks overlook the cognitive processes driving FM-based decisions, particularly in complex multi-agent FMware systems where outcomes can be non-deterministic or emergent (e.g., from interactions between agents or misinterpretation of prompts).
@@ V-B Observability
V-C Controlled Execution
Overview. Controlled execution in FMware development is essential for ensuring predictable, reliable, and efficient system behavior. However, as highlighted in Section IV, FMware presents a unique set of challenges due to its non-deterministic nature and the dynamic interactions between agents and models. Key recurrent issues include the lack of controlled execution mechanisms, which complicates the verification of fixes and limiting execution paths in FMware systems [82]. This lack of control severely limits debugging, reduces productivity, and can degrade system reliability, especially in multi-agent FMware where the same input might lead to divergent outputs across different executions.
@@ V-C Controlled Execution
A major issue is that the same input in FMware can lead to different execution paths with varying outputs due to the unpredictability of the underlying FMs and agents. Traditional techniques, which assume deterministic behavior, are not equipped for these variations, leading to flaky tests [10]. FMware often lacks repeatable execution, making it difficult to ensure consistent behavior after fixing issues. The absence of controlled execution frameworks hinders exploratory testing, preventing efficient identification of failure points and limiting performance optimizations. Without execution space restriction mechanisms like feature flags, development and maintenance become even more challenging. Additionally, the lack of detailed release notes and the disconnect between improvements and features make it hard to verify whether updates are effective. This problem is worse with cloud-hosted models, where developers have limited control over testing. Exploring and testing multiple execution paths in a structured way is essential for improving software quality and building user trust.
@@
– Repeatability. The framework should guarantee that the same input always produces the same execution flow, regardless of changes in external context or FM state. This could be achieved through execution snapshots and managing monosemantic units i.e., interpretable FM units that correspond to specific data patterns. By restoring snapshots and activating monosemantic units, developers can reproduce exact conditions that led to a bug or failure, facilitating debugging and fix verification. This approach is similar to deterministic replay systems used in distributed software, adapted for FMware’s dynamic and probabilistic nature [103].
@@
V-D Resource-Aware QA
Overview. In FMware production, managing high costs and resource demands is crucial, especially given the intensive requirements of FM operations. Traditional QA methods often overlook these needs, as FMware is hindered by low data efficiency, latency handling issues, high regression testing costs, and inefficient retry optimizations (refer Section IV). Unlike typical high-cost web services, FMs require significant computational power, have non-deterministic outputs, and need meticulous management of context and ability to quickly swap FMs. Resource-aware QA is essential for reducing the computational burden that FMware impose, particularly in large-scale environments where frequent calls to FMs can become prohibitively expensive.
Critical Analysis of the State-of-the-Practice. Current QA frameworks for FMware, adapted from traditional software engineering, fail to address FMware’s resource challenges. The problem is not just the absence of caching but the lack of FMware-native caching strategies that account for the dynamic nature of these systems. While solutions like LangChain [104] use traditional caching, the real challenge is deciding when to cache FM calls and ensuring cached responses remain relevant. FMware often updates databases and integrates real-time feedback, which makes cache invalidation critical. Without dynamic caching, systems risk re-running unnecessary queries or using outdated data, increasing the time and cost of regression testing.
Additionally, there is no systematic approach to optimize or prioritize the number of tests that must be run. Dependency-based test execution [105], effective in traditional software, has not been adapted for FMware. As a result, redundant tests waste computational resources, especially where FM calls incur significant costs in terms of latency and finances.
@@ V-D Resource-Aware QA
Our Vision. To address these challenges, we propose the development of a resource-aware QA framework specifically designed for FMware. This framework would incorporate a robust FMware-native caching mechanism that intelligently stores, updates and reuses FM query results, significantly reducing the need for repeated FM calls and maintaing cache integrity. By caching the results of expensive FM queries and reusing them across multiple tests, the system can reduce computational overhead while maintaining the accuracy of the testing process. This would be particularly valuable in scenarios where large volumes of tests are conducted against the same FM. In addition to caching, the framework should optimize test execution by dynamically prioritizing tests based on resource intensity. Grouping similar tests or those with expected outputs could minimize FM calls, lowering resource consumption. Test order optimization would also prioritize critical tests that are likely to surface defects, ensuring early issue detection. Native prompt compression techniques [106] should further reduce resource demands and improve latency.
@@
V-E Feedback Integration
Overview. Efficient feedback integration is vital for continuous improvement, optimization, and trustworthiness in production-ready FMware. However, as outlined in Section IV, two key issues often hinder this process: the lack of efficient feedback technology and cumbersome, error-prone memory management across different FMware systems. The absence of seamless feedback loops slows down FMware evolution and reduces adaptability, especially in complex real-world applications. Integrating feedback is further complicated by the need to manage user-specific and generalized knowledge in a scalable and non-disruptive way. As FMware becomes more widely adopted, a robust framework for feedback integration is necessary to ensure reliability, efficiency, and alignment with real-world demands.
Critical Analysis of the State-of-the-Practice. Feedback integration in FMware is underdeveloped compared to traditional software or machine learning systems, where feedback-driven optimization is well-established. The complexity of FMs and the unique characteristics of FMware complicate the application of these principles. A major issue is the lack of effective feedback mechanisms; most systems rely on ad-hoc methods requiring explicit user interactions, such as ratings [107], which suffice for demonstrations but are inadequate for production settings. Production environments demand passive, scalable, and continuous feedback collection to enable real-time insights. Without such mechanisms, developers miss essential opportunities for timely FMware refinement, hampering the correction of biases and inaccuracies and leading to slower adaptation cycles.
@@ V-E Feedback Integration
Additionally, the challenge of memory management, especially across different FMware systems, compounds the difficulties of feedback integration. Managing the context and knowledge required to incorporate feedback effectively is still error-prone, with systems struggling to maintain consistency across multiple feedback loops [108]. Without a robust system for synchronizing and managing feedback across different FMs and user contexts, the feedback process remains fragmented and unreliable.
@@
Knowledge Partitioning Framework. We propose developing a framework to manage different types of feedback by distinguishing between “outer knowledge” and “inner knowledge.” By automating the classification of feedback into these categories, FMware systems can ensure that general improvements are propagated across all users while maintaining user-specific optimizations where appropriate. The framework would also possess federated learning capabilities to leverage “inner knowledge” for global FMware improvement [109]. This partitioning is critical for scalable feedback integration, preventing overgeneralization of feedback while ensuring valuable information is not lost.
V-F Built-in Quality
Overview. Built-in quality is crucial for production-ready FMware to prevent costly rework and failures. As highlighted in Section IV, recurrent issues such as low domain coverage, poor data quality, inadequate prompt validation, and lack of guardrails for hallucinations stem from an over-reliance on FMs without properly structured knowledge and action spaces. Over-reliance on FMs elicits inefficiencies and unpredictable behavior. Addressing these concerns ensures FMware systems are robust, reliable, and scalable for real-world use, not just impressive demos.

### v2

@@ 4.11. Cross-cutting Issues that affect productionizing FMware
Lack of Automated Testing Capabilities. The lack of automated testing in FMware hinders efficient issue detection and continuous integration. Existing automated tools often miss the nuanced complexities in FMware, making them inadequate for production systems where manual testing is impractical due to scale. Without automation, development slows, increasing costs and the risk of undetected regressions (Abeysinghe and Circi, 2024; Hassan et al., 2024b).
@@
Lack of Assertion-Based Unit Tests. Writing assertion-based unit tests for FMware is difficult due to the non-deterministic nature of FMs, making expected outputs hard to define (Abeysinghe and Circi, 2024; sco, 2023). Traditional tests rely on fixed outcomes, but FMware’s varying outputs complicate this, especially in agent construction and memory management (sco, 2023). Despite these challenges, such tests are critical in production environments to detect regressions during refactoring or updates (Abeysinghe and Circi, 2024).
@@
Text-Based Evaluation Leads to Overestimation of Quality. Text-based evaluation in FMware often leads to an overestimation of quality by failing to capture deeper issues like relevance, accuracy, or alignment with business rules. Gao et al. (2024) highlight that current text-based Natural Language Generation (NLG) evaluations often miss key aspects such as context and factual correctness. Similarly, Hasanbeig et al. (2023) argue that relying solely on surface-level outputs can cause models to overfit to specific evaluation prompts, masking underlying performance issues.
@@
Ineffective and Inefficient AI-as-Judge Technologies. AI-as-Judge systems face inefficiency and ineffectiveness due to high costs and inherent biases. Thakur et al. (2024) highlight that models often struggle with hallucinations and inaccuracies, making their outputs unreliable in complex decision-making tasks. Additionally, the computational expense of using advanced models like GPT-4 makes them impractical for ongoing evaluations, especially when they fail to meet specific business needs (Thakur et al., 2024; Unknown, [n. d.]). These limitations, combined with the need for custom evaluation pipelines, reduce the feasibility of deploying such systems in production FMware.
@@
Difficulty navigating the regulatory and legal compliance minefield. Legal compliance and licensing complexities permeate all stages of the FMware lifecycle. FMs must align with the legal requirements of their deployment regions; for example, using ChatGPT-4-based FMware in China may not be viable due to local regulations (Hacker et al., 2023; Kelly, [n. d.]). Licensing FM data, including synthetic data, RLHF data, and user feedback, further complicates compliance. If a model generates synthetic data, the licensing of the data it was trained on affects the licensing of the generated data. Similarly, integrating user feedback (e.g., ratings or thumbs up/down) into FMware complicates matters if users later opt out, as removing their data from the FM and related components becomes difficult. Traditional Software Bills of Materials (SBOMs) ensure compliance for conventional software, but FMware’s complexity (encompassing data, models, and FM-generated data) goes beyond what traditional SBOMs can track.
5. Challenges
@@
Table 1. Challenges and Recurrent Issues in Developing Production-Ready FMware
@@
In this section, we present the results of our semi-structured thematic synthesis and outline the key challenges in productionizing FMware. It is important to note that the challenges presented in this version of the paper address only a subset of the identified recurrent issues, as shown in Table 1. Future versions will expand on this by incorporating additional challenges that cover all the recurrent issues across the FMware lifecycle.
@@
5.1. Testing
Overview. Poor testing practices present significant roadblocks in the transition from demo FMware to production-ready systems. As identified in Section 4, the lack of robust, assertion-based unit tests, the absence of automated testing, and the over-reliance on surface-level text-based evaluation (Abeysinghe and Circi, 2024; sco, 2023; Hassan et al., 2024b) often lead to overestimation of FMware quality. In real-world production settings, they result in unreliable outputs, latent defects, and increased regression costs. To address these issues, FMware require innovative testing mechanisms that account for the unique complexities of FMs and their non-deterministic behavior.
@@ 5.1. Testing
Moreover, the role of AI-as-judge technologies in testing pipelines adds another layer of complexity. Current AI judges, particularly those using large FMs like GPT-4, exhibit several critical flaws in evaluating correctness. These technologies often favor outputs based on superficial attributes, such as formatting or the length of responses rather than factual accuracy or alignment with real-world constraints (Thakur et al., 2024; Chen et al., 2024a). For instance, research shows that simply reversing the order of candidate responses can dramatically alter the AI’s judgment, resulting in false positives or negatives. The failure to maintain consistency in judging introduces significant risks when scaling FMware for production environments. Research has also points to a misalignment between AI decisions and human expectations. For example, Koo et al. (Koo et al., 2024) calculate the Rank-Biased Overlap (RBO) score to measure the agreement between human preferences and model evaluations in ranking-generated texts across 16 different LLMs (RBO varies from 0 to 1 and higher values indicate higher agreement). The average RBO was 0.44, with GPT-4 scoring 0.47. Another major issue arises from the associated cost and latency with AI-as-judge systems. Relying on large models like GPT-4 is computationally expensive and not viable for long-term or large-scale production-grade testing (Parnin et al., 2023).
@@
– Next-generation AI-judge framework. The development of such a framework is crucial for improving the reliability of evaluations. Unlike existing solutions, this framework would guide developers in crafting tailored prompts that align with specific business logic and domain constraints. These judges would be trained to focus not just on the superficial aspects of FM outputs (e.g., format, length) but on deeper attributes like factual accuracy, consistency with previous outputs, and compliance with application-specific requirements. To train the judges in a more structured, guided, flexible, evolvable, and cost-effective manner, we foresee the use of curriculum engineering (more details in Section 5.6). As a result, we would also obtain a more lightweight and efficient model, further contributing to cost-effectiveness.
5.2. Observability
Overview. Observability in FMware systems is crucial to ensure transparency, traceability, and reliability throughout their lifecycle. However, as outlined in Section 4, FMware presents unique challenges in observability due to its reliance on dynamic, non-deterministic components like FMs and autonomous agents. Recurrent issues, such as the complexity in determining the rationale behind system failures (i.e., whether stemming from limitations in prompts or FMs themselves) and the lack of FMware-native observability mechanisms (Hassan et al., 2024b), highlight the need for novel solutions. Unlike traditional software, FMware systems demand observability approaches that can capture both functional and cognitive aspects of AI agents and models.
Critical Analysis of the State-of-the-Practice. Current observability tools in FMware, while evolving, remain grounded in classical software observability practices. Tools like OpenLLMetry (tra, [n. d.]) and LangSmith (lan, [n. d.]) focus on low-level resource monitoring, capturing traces of FM calls, Vector DB interactions, and user prompts, along with metrics like latency and resource utilization. However, these frameworks overlook the cognitive processes driving FM-based decisions, particularly in complex multi-agent FMware systems where outcomes can be non-deterministic or emergent (e.g., from interactions between agents or misinterpretation of prompts).
@@ 5.2. Observability
5.3. Controlled Execution
Overview. Controlled execution in FMware development is essential for ensuring predictable, reliable, and efficient system behavior. However, as highlighted in Section 4, FMware presents a unique set of challenges due to its non-deterministic nature and the dynamic interactions between agents and models. Key recurrent issues include the lack of controlled execution mechanisms, which complicates the verification of fixes and limiting execution paths in FMware systems (Wu et al., 2024a). This lack of control severely limits debugging, reduces productivity, and can degrade system reliability, especially in multi-agent FMware where the same input might lead to divergent outputs across different executions.
@@ 5.3. Controlled Execution
A major issue is that the same input in FMware can lead to different execution paths with varying outputs due to the unpredictability of the underlying FMs and agents. Traditional techniques, which assume deterministic behavior, are not equipped for these variations, leading to flaky tests (Parnin et al., 2023). FMware often lacks repeatable execution, making it difficult to ensure consistent behavior after fixing issues. The absence of controlled execution frameworks hinders exploratory testing, preventing efficient identification of failure points and limiting performance optimizations. Without execution space restriction mechanisms like feature flags, development and maintenance become even more challenging. Additionally, the lack of detailed release notes and the disconnect between improvements and features make it hard to verify whether updates are effective. This problem is worse with cloud-hosted models, where developers have limited control over testing. Exploring and testing multiple execution paths in a structured way is essential for improving software quality and building user trust.
@@
– Repeatability. The framework should guarantee that the same input always produces the same execution flow, regardless of changes in external context or FM state. This could be achieved through execution snapshots and managing monosemantic units i.e., interpretable FM units that correspond to specific data patterns. By restoring snapshots and activating monosemantic units, developers can reproduce exact conditions that led to a bug or failure, facilitating debugging and fix verification. This approach is similar to deterministic replay systems used in distributed software, adapted for FMware’s dynamic and probabilistic nature (Alvaro and Quinn, 2024).
@@
5.4. Resource-Aware QA
Overview. In FMware production, managing high costs and resource demands is crucial, especially given the intensive requirements of FM operations. Traditional QA methods often overlook these needs, as FMware is hindered by low data efficiency, latency handling issues, high regression testing costs, and inefficient retry optimizations (refer Section 4). Unlike typical high-cost web services, FMs require significant computational power, have non-deterministic outputs, and need meticulous management of context and ability to quickly swap FMs. Resource-aware QA is essential for reducing the computational burden that FMware impose, particularly in large-scale environments where frequent calls to FMs can become prohibitively expensive.
Critical Analysis of the State-of-the-Practice. Current QA frameworks for FMware, adapted from traditional software engineering, fail to address FMware’s resource challenges. The problem is not just the absence of caching but the lack of FMware-native caching strategies that account for the dynamic nature of these systems. While solutions like LangChainuse traditional caching, the real challenge is deciding when to cache FM calls and ensuring cached responses remain relevant. FMware often updates databases and integrates real-time feedback, which makes cache invalidation critical. Without dynamic caching, systems risk re-running unnecessary queries or using outdated data, increasing the time and cost of regression testing.
Additionally, there is no systematic approach to optimize or prioritize the number of tests that must be run. Dependency-based test execution (Sharif et al., 2021), effective in traditional software, has not been adapted for FMware. As a result, redundant tests waste computational resources, especially where FM calls incur significant costs in terms of latency and finances.
@@ 5.4. Resource-Aware QA
Our Vision. To address these challenges, we propose the development of a resource-aware QA framework specifically designed for FMware. This framework would incorporate a robust FMware-native caching mechanism that intelligently stores, updates and reuses FM query results, significantly reducing the need for repeated FM calls and maintaing cache integrity. By caching the results of expensive FM queries and reusing them across multiple tests, the system can reduce computational overhead while maintaining the accuracy of the testing process. This would be particularly valuable in scenarios where large volumes of tests are conducted against the same FM. In addition to caching, the framework should optimize test execution by dynamically prioritizing tests based on resource intensity. Grouping similar tests or those with expected outputs could minimize FM calls, lowering resource consumption. Test order optimization would also prioritize critical tests that are likely to surface defects, ensuring early issue detection. Native prompt compression techniques (Wingate et al., 2022) should further reduce resource demands and improve latency.
@@
5.5. Feedback Integration
Overview. Efficient feedback integration is vital for continuous improvement, optimization, and trustworthiness in production-ready FMware. However, as outlined in Section 4, two key issues often hinder this process: the lack of efficient feedback technology and cumbersome, error-prone memory management across different FMware systems. The absence of seamless feedback loops slows down FMware evolution and reduces adaptability, especially in complex real-world applications. Integrating feedback is further complicated by the need to manage user-specific and generalized knowledge in a scalable and non-disruptive way. As FMware becomes more widely adopted, a robust framework for feedback integration is necessary to ensure reliability, efficiency, and alignment with real-world demands.
Critical Analysis of the State-of-the-Practice. Feedback integration in FMware is underdeveloped compared to traditional software or machine learning systems, where feedback-driven optimization is well-established. The complexity of FMs and the unique characteristics of FMware complicate the application of these principles. A major issue is the lack of effective feedback mechanisms; most systems rely on ad-hoc methods requiring explicit user interactions, such as ratings (Wang et al., 2021), which suffice for demonstrations but are inadequate for production settings. Production environments demand passive, scalable, and continuous feedback collection to enable real-time insights. Without such mechanisms, developers miss essential opportunities for timely FMware refinement, hampering the correction of biases and inaccuracies and leading to slower adaptation cycles.
@@ 5.5. Feedback Integration
Additionally, the challenge of memory management, especially across different FMware systems, compounds the difficulties of feedback integration. Managing the context and knowledge required to incorporate feedback effectively is still error-prone, with systems struggling to maintain consistency across multiple feedback loops (Peng et al., 2023). Without a robust system for synchronizing and managing feedback across different FMs and user contexts, the feedback process remains fragmented and unreliable.
@@
Knowledge Partitioning Framework. We propose developing a framework to manage different types of feedback by distinguishing between “outer knowledge” and “inner knowledge.” By automating the classification of feedback into these categories, FMware systems can ensure that general improvements are propagated across all users while maintaining user-specific optimizations where appropriate. The framework would also possess federated learning capabilities to leverage “inner knowledge” for global FMware improvement (Yamamoto et al., 2023). This partitioning is critical for scalable feedback integration, preventing overgeneralization of feedback while ensuring valuable information is not lost.
5.6. Built-in Quality
Overview. Built-in quality is crucial for production-ready FMware to prevent costly rework and failures. As highlighted in Section 4, recurrent issues such as low domain coverage, poor data quality, inadequate prompt validation, and lack of guardrails for hallucinations stem from an over-reliance on FMs without properly structured knowledge and action spaces. Over-reliance on FMs elicits inefficiencies and unpredictable behavior. Addressing these concerns ensures FMware systems are robust, reliable, and scalable for real-world use, not just impressive demos.